Calculate the commission based on the sales

data in column B and insert the calculated commission into column C using the rules in the

table below.

Use a nested IF statement to solve this problem (Can you write the IF Statement? Thank you! I also formatted it to how it would look in an excel doc).

Column A                            Column B                  Column C

A1 SALESPERSON B1 SALES C1 COMMISSION

A2 Anderson B2 1255 C2

A3 Kwan B3 235 C3

A4 Kean    B4 750 C4

A5 Solomon B5 450 C5

For sales less than $400. The commission is 7%

For sales from $401.to $700. The commission is $750. plus 10% commission

For sales from $701. To 1000. The commission is 12.5%

For sales from $1001. and higher the commission is 16%
Answer #1

Dear Student ,

As per the requirement submitted above , kindly find the below solution.

Excel IF Formula :

=IF(B2<400,B2*7/100,IF(( AND(B2>=401,B2<700)),750+B2*10/100,IF(( AND(B2>=701,B2<1000)),B2*12.5/100,IF(B2>=1001,B2*16/100,0))))

Below is the screen in excel :

NOTE : PLEASE FEEL FREE TO PROVIDE FEEDBACK ABOUT THE SOLUTION.